OpenStudy (anonymous):

Which of the following is the equation of the axis of symmetry of the quadratic equation y = 2x2 + 24x + 62?

OpenStudy (anonymous):

ppp please help me, I have no idea how to do this

OpenStudy (anonymous):

Your equation is in the form y=ax^2+bx+c 2=a 24=b 62=c formula for axis of symmetry is x=-b/2a x=-24/(2*2) x=?

OpenStudy (anonymous):

x = –6 x = 6 x = –10 x = 10 these be the possible answers...

OpenStudy (anonymous):

And if you do have the time, i would very much appreciate an explanation to the problem ...... Describe the differences between the graph of y = –3(x + 7)2 – 10 and the standard position graph of y = x2.

OpenStudy (anonymous):

Did you read my post? All thats left to do is -24/(2*2)

OpenStudy (anonymous):

so -6 yea?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

thank you kind sir, I owe you one ;)